# FERGI: Automatic Scoring of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction

### v3

@@ 1 Introduction
However, the scalability of human feedback collection has still been limited by its reliance on manual annotation. Researchers have used the CLIP score [59], Aesthetic score [64], BLIP score [43], ImageReward score [79], PickScore [35], and HPS v2 score [77] to automatically score human preferences of text-to-image generation based on the input text prompts and the generated images. In this study, we develop and test a system to automatically score user preferences from their spontaneous facial expression reaction to the generated images, which is complementary to the models based only on the texts and images. This automatic scoring requires zero additional effort from the real users of text-to-image generators.
@@ 2 Related Work
Text-to-Image Generation and Evaluation. Various models have been developed for text-to-image generation, including Generative Adversarial Networks (GANs) [25, 32, 5, 31, 15, 16], Variational autoencoders (VAEs) [34, 10, 68], flow-based models [17, 18, 33], autoregressive models (ARMs) [9, 11, 71, 72, 84], and diffusion models (DMs) [66, 28, 67]. More recent advancements in DMs [54, 62, 61, 58, 63, 14] have achieved great success in generating high-quality images and attracted widespread public attention. As traditional evaluation metrics like Inception Score (IS) [3], Fréchet Inception Distance (FID) [27], and CLIP score [59] failed to comprehensively capture human preferences for text-to-image generation, researchers have proposed to specifically train human preference scoring models based on large-scale human feedback datasets, such as Human Preference Score (HPS) [78], ImageReward [79], PickScore [35], and Human Preference Score v2 (HPS v2) [77], and demonstrated the effectiveness of using them to improve the text-to-image generative models [79, 35, 77, 24].
@@
Facial Expression Recognition and Applications. Automatic facial expression recognition has advanced rapidly in recent years [44]. As emotional responses to image generation can be more complicated than single categories of the most widely-researched seven basic emotions [20, 19, 50], we decided to analyze the facial expression reactions directly in terms of muscle movements, as defined by AUs in FACS [21]. AU detection and estimation has attracted increasing interest [29, 65, 88, 87, 12, 45, 46, 73, 36, 37, 89, 49, 69, 70] and been used for facial emotion recognition [81, 83, 82, 80]. For example, researchers have shown that the activation of the corrugator (activated in AU4) and zygomaticus (activated in AU12) are associated with negative and positive emotions respectively and are the most investigated muscle activations in studying affect and emotion [76]. Also, the activity of the corrugator supercilii (activated in AU4) is positively associated with amygdala and negatively associated with ventromedial prefrontal cortex activity [26] and increases with negative stimuli, such as negative images [7, 6, 40, 39]. For applications, [4, 75] showed the effectiveness of using facial expression to infer/analyze human preferences.
@@ 3.1 Participants
39 participants were recruited from the SONA system of University of California San Diego (UCSD) and completed the study asynchronously on personal computers. 6 participants were removed from the dataset for various reasons, including failure to participate, failure in video recording, and not permitting sharing of their video recordings for research. This resulted in a dataset of 33 participants that will be made available for research. The collection and usage of human participant data in the dataset were approved by the Institutional Review Board (IRB) of UCSD.
@@ 3.2 Data Collection Procedure
Each participant completed multiple sessions of data collection. In each session, the participant drafted one text prompt and viewed 5 images generated from the input text prompt. A flow chart of the procedure of each session is shown in Fig. 1. Details are explained below.
Figure 1: Procedure of one session in data collection.
@@ 3.2.1 Prompt Creation
There are two types of sessions, structured input and free-form input. In both types, the participant can freely draft any text prompt as long as it is not NSFW. The only difference between the two types lies in how the input text prompt is created. In a structured input session, the participant fills out a form to separately specify different elements desired in the images to generate: animate objects, inanimate objects, interactional relations (between two objects), positional relations (between two objects), location, style (of the image), and keywords. Based on the contents of the form, a text prompt is then automatically generated by the Large Language Model (LLM) OpenAI gpt3.5-turbo. The participant can freely adjust the generated prompt as the final input text prompt. In a free-form input, the participant directly enters the entire final input text prompt.
@@ 3.2.2 Recording
After the input text prompt is finalized, the participant is directed to a webcam preview to confirm that the webcam captures their face appropriately. The webcam is then used to record throughout the rest of the session. The configuration of the raw videos varied based on the participant’s device and browser. Though, recorded videos of most participants have a resolution of $640\times 480$, and all recorded videos were standardized to 30 FPS afterwards.
@@ 3.2.3 Image Presentation, Annotation, and Ranking
In each session, a total of 5 images generated by SD v1.4 [62] from the input text prompt are presented and annotated sequentially and ranked afterwards (present image 1 $\rightarrow$ annotate image 1 $\rightarrow$ $\dots$ $\rightarrow$ present image 5 $\rightarrow$ annotate image 5 $\rightarrow$ rank 5 images). The image presentation and annotation followed the webcam preview. The detailed procedures for the presentation and annotation of each generated image are as follows:
  • Baseline (5 seconds): The website displays the “preparing image” status text at the center of the page along with the input text prompt. No specific facial-expression-eliciting event is expected to happen during this period.
@@
  • Annotate (no time restriction): The website displays the prompt, the generated image (moved to the right side), and the annotation survey. The participant annotates the image manually by filling out the survey during this interval. The survey was adapted from the survey designed for ImageReward [79] and includes the following questions: important elements not reflected in the image (free-response question), overall rating (on a scale of 1 (worst) to 7 (best)), image-text alignment rating (on a scale of 1 to 7), fidelity rating (on a scale of 1 to 7), issues of the image [^1] For the first 7 out of 33 participants (or 8 out of the total 39 participants), there was a bug (later fixed) in checkbox selection for answering the questions regarding “issues of the image” that made the recorded answers unreliable, so the data regarding their answers for this question should not be used in analysis. (multi-choice question), and emotions felt when seeing the image (multi-choice question).
@@ 4 AU Model Training
We analyze the facial expression reaction of the users by estimating the activation of their AUs. We trained our AU estimation model on the DISFA [53, 52] and DISFA+ [51] datasets. They are among the largest, most popular, and most cited AU datasets with intensity annotations. In these datasets, each frame is manually annotated by a human expert with intensities on a scale of 0 to 5, for activation of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop). (See Fig. S10 for a visual reference guide for the analyzed AUs.) Video frames are first preprocessed with face detection [48], face alignment [8], and a combination of histogram equalization and linear mapping [38] before being fed into the model for training, testing, or inference. For model training, we used the neural network IR-50 [13] pre-trained on Glink360k [1] and fine-tuned it on the DISFA and DISFA+ datasets [53, 51]. A single network learned the estimation of all labeled AUs with both regression and ordinal classification [55]. More details of the training of the AU model is elaborated in the Appendix / Supplementary Materials.
The comparison between our AU estimation model’s performance on the DISFA dataset and other state-of-the-art models, including OpenFace 2.0 [2], CCNN-IT [73], 2DC [47], SCC-Heatmap [23], and iARL [65] is shown in Tab. 1. Our model’s Pearson’s $r$ (higher is better) is close to that of OpenFace 2.0; our model’s ICC(3,1) (higher is better) is the best among all models. Although our model’s MAE (lower is better) is less competitive among all models, we believe that ICC(3,1) is a better metric for evaluating the performance of AU estimation models on the DISFA dataset considering the high imbalance of the dataset.
@@
AU 1 2 4 5 6 9 12 15 17 20 25 26 Avg.
Pearson’s $r$ (higher is better)
OpenFace 2.0 .64 .50 .70 .67 .59 .54 .85 .39 .49 .22 .85 .67 .59
Our Model .53 .46 .74 .57 .56 .60 .85 .42 .41 .21 .94 .66 .58
ICC(3,1) (higher is better)
CCNN-IT .18 .15 .61 .07 .65 .55 .82 .44 .37 .28 .77 .54 .45
2DC .70 .55 .69 .05 .59 .57 .88 .32 .10 .08 .90 .50 .50
SCC-Heatmap .73 .44 .74 .06 .27 .51 .71 .04 .37 .04 .94 .78 .47
iARL .13 .36 .68 .22 .56 .36 .86 .52 .37 .12 .96 .60 .48
Our Model .53 .46 .74 .56 .55 .57 .85 .41 .40 .20 .93 .66 .57
MAE (lower is better)
CCNN-IT 87 .63 .86 .26 .73 .57 .55 .38 .57 .45 .81 .64 .61
SCC-Heatmap .16 .16 .27 .03 .25 .13 .32 .15 .20 .09 .30 .32 .20
iARL .30 .31 .52 .04 .36 .30 .31 .05 .33 .08 .29 .26 .26
Our Model .37 .39 .50 .14 .37 .21 .33 .24 .47 .23 .32 .42 .33
@@ 5 Facial Feature Extraction
In this section, we elaborate the procedure of extracting facial features from the video clips associated with each generated image based on the trained AU recognition model. Similar procedures were used to extract features from baseline clips and reaction clips. Further details are in the Appendix / Supplementary Materials.
@@ 5.2 AU Activation Value
    $\bar{\hat{y}}_{i,\mathrm{reg}}^{(k)}=\frac{1}{3}\Sigma_{k^{\prime}=k}^{k+2}\hat{y}_{i,\mathrm{reg}}^{(k^{\prime})}$ (1)
@@ 6.1 Statistical Analysis
(a) surprised
(b) disgusted
(c) amused
(d) scared
(e) disappointed
(f) satisfied
Figure 3: Reported emotions highly correlated with AU activation values. The distributions of the activation values of multiple AUs for images eliciting different emotions of the participants as reported in answers to the question “Did you feel any of the following emotions when you saw the image?”. Each subfigure shows results for a different emotion (indicated in the captions) and contains multiple pairs of boxplots representing the results for different AUs. The indices of the AUs are along the x-axis, while the y-axis represents the AU activation values in the reaction clip of each image. The parenthesized numbers in the legends give the number of responses of that type. In each pair of boxplots, the boxplot on the left/right side represents the AU activation values for images that did/didn’t elicit the corresponding emotion. In each boxplot, the bottom/top of the box represents the first/third quartile (25th/75th percentile) of the AU activation values, and the line in the middle of the box represents the median.
@@
  • Overall Ratings. The participant’s overall ratings (on a scale of 1 to 7) of the generated images have a significant positive correlation with the activation values of AU2 and AU12 and a significant negative correlation with the activation values of AU4 and AU9 as shown in Figs. 2(f), 2(a), 2(b) and 2(g) (Spearman correlation: $\rho\approx 0.09$ and $p<5\times 10^{-5}$ for AU2, $\rho\approx-0.18$ and $p<5\times 10^{-17}$ for AU4, $\rho\approx-0.09$ and $p<5\times 10^{-5}$ for AU9, $\rho\approx 0.08$ and $p<5\times 10^{-4}$ for AU12).
@@
    • Surprised. “Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig. 3(a) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
    • Disgusted. “Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig. 3(b) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
    • Amused. “Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig. 3(c) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
    • Scared. “Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig. 3(d) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
    • Disappointed. “Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Fig. 3(e) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
    • Satisfied. “Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Fig. 3(f) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).

### v4

@@ 1 Introduction
However, the scalability of human feedback collection has still been limited by its reliance on manual annotation. Researchers have used the CLIP score [60], Aesthetic score [65], BLIP score [43], ImageReward score [80], PickScore [35], and HPS v2 score [78] to automatically score human preferences of text-to-image generation based on the input text prompts and the generated images. In this study, we develop and test a system to automatically score user preferences from their spontaneous facial expression reaction to the generated images, which is complementary to the models based only on the texts and images. This automatic scoring requires zero additional effort from the real users of text-to-image generative models.
@@ 2 Related Work
Text-to-Image Generation and Evaluation. Various models have been developed for text-to-image generation, including Generative Adversarial Networks (GANs) [25, 32, 4, 31, 15, 16], Variational autoencoders (VAEs) [34, 10, 69], flow-based models [17, 18, 33], autoregressive models (ARMs) [9, 11, 72, 73, 86], and diffusion models (DMs) [67, 28, 68]. More recent advancements in DMs [54, 63, 62, 59, 64, 14] have achieved great success in generating high-quality images and attracted widespread public attention. As traditional evaluation metrics like Inception Score (IS) [2], Fréchet Inception Distance (FID) [27], and CLIP score [60] failed to comprehensively capture human preferences for text-to-image generation, researchers have proposed to specifically train human preference scoring models based on large-scale human feedback datasets, such as Human Preference Score (HPS) [79], ImageReward [80], PickScore [35], and Human Preference Score v2 (HPS v2) [78], and demonstrated their effectiveness to improve text-to-image generative models [80, 35, 78, 24].
@@
Facial Expression Recognition and Applications. Automatic facial expression recognition has advanced rapidly in recent years [44]. As emotional responses to image generation can be more complicated than single categories of the most widely-researched seven basic emotions [20, 19, 50], we decided to analyze the facial expression reactions directly in terms of muscle movements, as defined by AUs in FACS [21]. AU detection and estimation has attracted increasing interest [29, 66, 91, 90, 12, 45, 46, 74, 36, 37, 92, 49, 70, 71] and been used for facial emotion recognition [83, 85, 84, 82, 81]. For example, researchers have shown that the activation of the corrugator (activated in AU4) and zygomaticus (activated in AU12) are associated with negative and positive emotions respectively and are the most investigated muscle activations in studying affect and emotion [77]. Also, the activity of the corrugator supercilii (activated in AU4) is positively associated with amygdala and negatively associated with ventromedial prefrontal cortex activity [26] and increases with negative stimuli, such as negative images [7, 6, 40, 39]. For applications, [3, 76] showed the effectiveness of using facial expression to infer/analyze human preferences.
@@ 3.1 Participants
39 participants were recruited from the SONA system of University of California San Diego (UCSD) and completed the study asynchronously on personal computers. 6 participants were removed from the dataset for various reasons, including failure to participate, failure in video recording, and not allowing sharing of their video for research. Data collection and use were approved by the Institutional Review Board (IRB) of UCSD.
@@ 3.2 Data Collection Procedure
Each participant completed multiple sessions of data collection. In each session, the participant drafted one text prompt and viewed 5 images generated from the input text prompt. A flow chart of the procedure of each session is shown in Figure 1. Details are explained below.
Figure 1: Procedure of one session in data collection. Firstly, the participant creates an input text prompt. Secondly, the participant is directed to a webcam preview to confirm that the webcam captures their face appropriately; at the same time, 5 images are generated from the text prompt using Stable Diffusion v1.4. Then, for each image, the participant goes through a 5-second baseline phase, a 5-second image presentation phase, and an image annotation phase with no time restriction. (During the baseline phase, the participant’s baseline facial expression is recorded, and during the image presentation phase, the participant’s facial expression reaction to the generated image is recorded.) Finally, after all 5 images have been presented and annotated, the participant ranks the 5 generated images from best to worst. A copy of this figure with additional references to the screenshots of each stage is in the supplemental material.
@@ 3.2.1 Prompt Creation
There are two types of sessions, structured input and free-form input. In both types, the participant can freely draft any text prompt as long as it is not NSFW. The only difference lies in how the input text prompt is created. In a structured input session, the participant fills out a form to separately specify different elements desired in the generated images: animate objects, inanimate objects, interactional relations (between two objects), positional relations (between two objects), location, style (of the image), and keywords. Based on the contents of the form, an initial text prompt is automatically generated by the Large Language Model (LLM) OpenAI GPT-3.5 Turbo which the participant can freely adjust to obtain the final input text prompt. In a free-form input, the participant directly enters the entire final input text prompt.
@@ 3.2.2 Recording
After the input text prompt is finalized, the participant is directed to a webcam preview to confirm that the webcam captures their face appropriately. The webcam is used to record throughout the rest of the session. The configuration of the raw videos varied based on the participant’s device and browser. However, recorded videos of most participants have a resolution of $640\times 480$, and all recorded videos were standardized to 30 FPS afterwards.
@@ 3.2.3 Image Presentation, Annotation, and Ranking
In each session, a total of 5 images generated by SD v1.4 [63] from the input text prompt are presented and annotated sequentially and ranked afterwards (present image 1 $\rightarrow$ annotate image 1 $\rightarrow$ $\dots$ $\rightarrow$ present image 5 $\rightarrow$ annotate image 5 $\rightarrow$ rank 5 images). The image presentation and annotation followed the webcam preview. The details are as follows:
  • Baseline (5 seconds): The website displays the “preparing image” status text at the center of the page along with the input text prompt. No specific facial-expression-eliciting event is expected during this period.
@@
  • Annotate (no time restriction): The website displays the prompt, the generated image (moved to the right side), and the annotation survey. The participant annotates the image manually by filling out the survey during this interval. The survey was adapted from the survey designed for ImageReward [80] and includes the following questions: important elements not reflected in the image (free-response question), overall rating (on a scale of 1 (worst) to 7 (best)), image-text alignment rating (on a scale of 1 to 7), fidelity rating (on a scale of 1 to 7), issues of the image [^1] For the first 7 out of 33 participants (or 8 out of the total 39 participants), there was a bug (later fixed) in checkbox selection for answering the questions regarding “issues of the image” that made the recorded answers unreliable, so answers for this question should not be used in analysis. (multi-answer question), and emotions felt when seeing the image (multi-answer question).
@@ 4 AU Model Training
We analyze the facial expression reaction of the users by estimating their AU activations. We trained our AU estimation model on the DISFA [53, 52] and DISFA+ [51] datasets. They are among the largest, most popular, and most cited AU datasets with intensity annotations. In these datasets, each frame is manually annotated by a human expert with intensities on a scale of 0 to 5, for activation of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop). (See the supplemental material for a visual reference guide for the analyzed AUs.) Video frames are first preprocessed with face detection [48], face alignment [8], and a combination of histogram equalization and linear mapping [38] before being fed into the model for training, testing, or inference. For model training, we used the neural network IR-50 [13] pre-trained on Glink360k [1] and fine-tuned it on the DISFA and DISFA+ datasets [53, 51]. A single network learned the estimation of all labeled AUs with both regression and ordinal classification [56]. More details of the training of the AU model are elaborated in the supplemental material.
The comparison between our AU estimation model’s performance on the DISFA dataset and other state-of-the-art models is shown in Table 1. Our model’s ICC(3,1) (higher is better) is the best among all models. Although our model’s MAE (lower is better) is less competitive among all models, we believe that ICC(3,1) is a better metric for evaluating the performance of AU estimation models on the DISFA dataset considering the high imbalance of the dataset. Also, ICC(3,1) is specifically more important in our application here because we care about the change of facial expression in reaction to the generated images instead of its absolute intensity.
@@
Metric Method AU Average
1 2 4 5 6 9 12 15 17 20 25 26
ICC(3,1$\uparrow$ CCNN-IT [74] .18 .15 .61 .07 .65 .55 .82 .44 .37 .28 .77 .54 .45
2DC [47] .70 .55 .69 .05 .59 .57 .88 .32 .10 .08 .90 .50 .50
SCC-Heatmap [23] .73 .44 .74 .06 .27 .51 .71 .04 .37 .04 .94 .78 .47
iARL [66] .13 .36 .68 .22 .56 .36 .86 .52 .37 .12 .96 .60 .48
[…]
SCC-Heatmap [23] .16 .16 .27 .03 .25 .13 .32 .15 .20 .09 .30 .32 .20
iARL [66] .30 .31 .52 .04 .36 .30 .31 .05 .33 .08 .29 .26 .26
[…]
@@ 5 Facial Feature Extraction
In this section, we elaborate the procedure of extracting facial features from the video clips associated with each generated image based on the trained AU recognition model. Similar procedures were used to extract features from baseline clips and reaction clips. Further details are in the supplemental material.
@@ 5.2 AU Activation Value
    $\bar{\hat{y}}_{i,\mathrm{reg}}^{(k)}=\frac{1}{3}\sum_{k^{\prime}=k}^{k+2}\hat{y}_{i,\mathrm{reg}}^{(k^{\prime})}$ (1)
@@ 6.1 Statistical Analysis
(a) disappointed
(b) satisfied
(c) surprised
(d) disgusted
(e) amused
(f) scared
Figure 3: Reported emotions highly correlated with AU activation values. The distributions of the activation values of multiple AUs for images eliciting different emotions of the participants as reported in answers to the question “Did you feel any of the following emotions when you saw the image?” Each subfigure shows the results for a different emotion (indicated in the subcaptions) and contains multiple pairs of boxplots representing the results for different AUs. The x-axis represents the indices of the AUs while the y-axis represents the AU activation values in the reaction clip of each image. In each pair of boxplots, the boxplot on the left/right side represents the AU activation values for images that did/didn’t elicit the corresponding emotion. In each boxplot, the bottom of the box represents the first quartile (25th percentile) of the AU activation values, the top of the box represents the third quartile (75th percentile), and the line in the middle of the box represents the median. The numbers of responses of each type for each subfigure are as follows. (a) Yes (592); No (1625). (b) Yes (724); No (1493). (c) Yes (352); No (1865). (d) Yes (124); No (2093). (e) Yes (368); No (1849). (f) Yes (194); No (2023).
@@
  • Overall Ratings. The participant’s overall ratings (on a scale of 1 to 7) of the generated images have a significant positive correlation with the activation values of AU2 and AU12 and a significant negative correlation with the activation values of AU4 and AU9 as shown in Figures 2(f), 2(a), 2(b) and 2(g) (Spearman correlation: $\rho\approx 0.09$ and $p<5\times 10^{-5}$ for AU2, $\rho\approx-0.18$ and $p<5\times 10^{-17}$ for AU4, $\rho\approx-0.09$ and $p<5\times 10^{-5}$ for AU9, $\rho\approx 0.08$ and $p<5\times 10^{-4}$ for AU12).
@@
    • Disappointed. “Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Figure 3(a) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
    • Satisfied. “Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Figure 3(b) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
    • Surprised. “Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Figure 3(c) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
    • Disgusted. “Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Figure 3(d) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
    • Amused. “Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Figure 3(e) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
    • Scared. “Scary” images are associated with higher activation values of AU4 and AU5 as shown in Figure 3(f) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).